what is this problem for the i need help please. x divided by 3 minus 6 equal x divided by 9?
Answer by mmm4444bot(95)

Hello There:
If I interpret your equation correctly, it is:
x/3 - 6 = x/9
Let's start by clearing the fractions. We do this by multiplying both sides of the equation the the least common multiple of the denominators. In this case, it is 9.
9*(x/3) - 9*6 = 9*(x/9)
3*x - 54 = x
Now we want to get all of the x-terms on the left side of the equals sign, and we want the constant on the right side. Subtract x from both sides; then add 54 to both sides.
3*x - x - 54 = x - x
3*x - x - 54 = 0
3*x - x + 54 = 0 + 54
3*x - x = 54
Combine the two x-terms on the left.
2*x = 54
The last step is to divide both sides by 2 to isolate x by itself.
(2/2)*x = 54/2
x = 27
~ Mark